Io, a satellite of Jupiter, has an orbital period
Answer #1

We know that the centripetal Force of the satellite is equal the Gravitational Force due to The Jupiter.

mV^2/r =GMm/r^2

M=V^2.r/G. ...1

M is mass of the Jupiter.

Here V=2\pir/T put V in equation 1

M=(2\pir)^2r/T^2G

M=4\pi^2r^3/T^2G. ..2

Here r=4.22×10^8 m , T =1.77days =152928 s G=6.674×10^(-11) N.m^2/kg^2

Put these values in equation 2

M=4\pi^2(4.22×10^8)^3/(152928)^2×6.674×10^(-11)

M=1.90080×10^(-8)×10^35

M=1.90×10^27 kg